Question 1 (1 point) 0.5E, the total damage cost of 20 units of For a marginal damage function equal to: MD emissions is __.- not possible to tell with the information given $100 $200 $10 Question 2 (1 point) Consider the following graph. What is the socially-efficient level of emissions? The equation for the marginal damages function is: MD 200 E MDC 900 200 MAC 1200 Emissions Question 3 (1 point) Suppose MAC1 100- 10E, MAC2 50 10E and MD 30E What is the socially-efficient level of emissions Question 4 (1 point) There are three polluting sources with the following functions: MAC1 100 4E1, MAC2 80 -2E2 and MAC3 60-0.5E3 With no regulations in place, what is likely to be the total amount of pollution emitted by all 3 sources?
Answer #1

(Question 1) Option (3)

When E = 20,

MD = 0.5 x 20 = $10

Total damage cost = MD x E = $10 x 20 = $200

(Question 2)

Equation of MAC curve: MAC = a - bE

When E = 0, MAC = 900 (Vertical intercept). So

900 = a

When E = 1200, MAC = 0 (Horizontal intercept). So

0 = a - 1200E

1200E = a = 900

E = 0.75

Therefore,

MAC = 900 - 0.75E

Equating with MD,

200 + E = 900 - 0.75E

1.75E = 700

E = 400

NOTE: As per Answering Policy, 1st 2 questions are answered.